OpenStudy (anitas.):

plz check my answer to this equation!!!-----> (x^3)^2 · x^-2

OpenStudy (hossamhoussien):

\[(x^3)^2 \times x ^{-2}\] Is that the equation??? And where is your answer?

OpenStudy (anitas.):

(x^3^)2 · x^-2 = x^3*2 · x^-2 = x^6 · x^-2 = x^6+(-2) = x^4

OpenStudy (anitas.):

That is my work^ and i need someone to check it plz

OpenStudy (hossamhoussien):

Yea it's correct ;)

OpenStudy (anitas.):

ok thank you very much!!! : ]

OpenStudy (hossamhoussien):

yw